# No Need to Look Back: An Efficient and Scalable Approach for Temporal Network Representation Learning

### v1

@@ Appendix D Dataset Description
  • GDELT [^4] https://github.com/tedzhouhk/TGL is a Temporal Knowledge Graph (TKG) generated from the GDELT Event Database (Leetaru & Schrodt, 2013) and adapted by TGL (Zhou et al., 2022) for richer node and edge features. It is a large-scale dataset with more than 100M links. It records events from news and articles in over 100 languages every 15 minutes. The nodes represents actors and the edges represent events that happen between a pair of actors. The node features represent the CAMEO codes of the actors and the link features represent the CAMEO codes of the events. The labels are the countries where the actors were located when the events happen.
@@
    MAG [^5] https://github.com/tedzhouhk/TGL is a homogeneous sub-graph of the heterogeneous MAG240M graph in OGB-LSC (Hu et al., 2021) extracted by TGL. It is another large-scale dataset with more than 100 million nodes and 1.3 billion links. It is a paper-paper citation network where each node in MAG represents one academic paper. The temporal edges represent citation of one paper to another with timestamp representing the year when the paper is published. The node features are the embeddings of the abstract of the paper generated with RoBERTa (Liu et al., 2019). The node labels are the arXiv subject areas.
@@
  • Ubuntu [^6] https://snap.stanford.edu/data/sx-askubuntu.html, is a dataset that records the events on a stack exchange web site called Ask Ubuntu. [^7] http://askubuntu.com/ The nodes represent users and there are three different types of edges, (1) user $u$ answering user $v$’s question, (2) user $u$ commenting on user $v$’s question, and (3) user $w$ commenting on user $u$’s answer.
@@
  • Wiki-talk [^8] https://snap.stanford.edu/data/wiki-talk-temporal.html is dataset that represents the edit events on Wikipedia user talk pages. The dataset spans approximately 5 years so it accumulates a large number of nodes and edges. This is a large-scale dataset with more than 1M nodes.
@@
Table 8: Fixed hyperparameters for all methods including NLB. GDELT and MAG use smaller $s$ because of CPU and GPU size limits.
@@ Appendix E Baselines, hyperparameters and the experiment setup
DySAT (Sankar et al., 2020) is a snapshot-based approach that aggregates features from local node neighborhoods in each snapshot through self-attention and uses RNNs to combine the node embeddings from different snapshots. We use 3 snapshots for all datasets.
@@
TGAT (Xu et al., 2020) adapts GAT (Veličković et al., 2018) for static graphs to dynamic graphs. It aggregates messages from dynamic neighbors via an attention mechanism. While it proposed recent sampling, its implementation is significantly inefficient because it has to calculate the recency weights online before sampling. Thus, we only consider uniform sampling and truncation. We use 2 attention heads and 100 hidden dimensions. We only consider aggregation from the first-hop neighbors because aggregating second-hop neighbors introduces even more severe latency for sampling. We also do not observe significant improvement in prediction performance using second-hop neighbors.
@@
TGN (Rossi et al., 2020) is a recently proposed SOTA temporal graph representation learning approach. It keeps track of a memory state for each node and update with new interactions. We train TGN with 100 dimensions for each of memory module, time feature and node embedding. We also only consider sampling the first-hop neighbors because of computational efficiency.
@@
APAN (Wang et al., 2021c) is a recent approach designed for low inference latency. It prepares the neighborhood information forward with graph propagation instead of aggregation. While it inspires our approach, it does not support recent sampling and has shown to consume significant CPU memory while performing subpar to SOTA methods. The reason is it stores the messages for each node in CPU which consumes massive CPU memory. We train APAN with 100 dimensions for the node embeddings and 10 dimensions for the mailbox.
@@
TGL (Zhou et al., 2022) is a very recent framework that support efficient and scalable temporal graph representation learning. It proposes temporal neighbor sampling with multi-core CPUs in parallel. Overall, it is shown to achieve on average 13$\times$ speedup for training and 173$\times$ speedup on neighbor sampling. For all of the baselines above, we evaluate them using the implementation by TGL. For the GDELT and the MAG datasets, since the node and edge features do not fit in our GPU, we set the all_on_gpu flag to false. For all other datasets, we set it to true. The source code is provided here. [^9] https://github.com/tedzhouhk/TGL
@@
NAT (Luo & Li, 2022) is also a very recent work that achieves SOTA link prediction performance for dynamic graphs. It proposed dictionary-type node representations to replace the traditional single-vector node representations which keep track of historical neighborhood information and avoid the expensive backward neighbor sampling. However, it relies on the construction of joint neighborhood features which only work with node pairs and cannot be easily generalized to node classification. Therefore, we evaluate it on link prediction only. We also only use the first-hop neighborhood. Other than that, we follow the hyperparameter settings documented in their paper.
@@
NLB. For both NLB-node and NLB-edge, we tune the node representation and the node status dimensions between 50 to 100. We use two attention heads and we tune $\alpha$ and dropout between 0 to 1 with grid search.
@@ E.1 Inductive evaluation of NLB
Our evaluation pipeline for inductive learning is similar to NAT (Luo & Li, 2022) and different from other baselines. For backward sampling methods such as TGN (Rossi et al., 2020) and TGAT (Xu et al., 2020), when they do inductive evaluations, they can access the entire training and evaluation data for aggregation, including events that are masked for inductive test. However, NLB’s forward sampling prepares the down-sampled neighbors forward for future use and ignores the masked events during the training phase. Thus, by the end of the training, even all historical events become accessible, NLB cannot leverage them. Therefore, to ensure a fair comparison, after training, NLB processes the full train and validation data with all nodes unmasked to gather the down-sampled temporal neighbors from the complete set, and then processes the test data. Note that in this last pass over the full train and validation data, we only perform the forward sampling and do not perform training anymore.
@@ E.2 Node classification preprocessing
Appendix F TGL Node Classification for MAG
We fail to reproduce the node classification scores for MAG over all of the baselines implemented in TGL (Zhou et al., 2022). Our hypothesis is as follows. TGL has reported the scores for MAG based on an implementation with an issue. For the backtracking and sampling of temporal neighbors to learn representation at time $t$, TGL considers temporal neighbors of the same timestamp $t$ which may cause information leakage. We noticed this issue in their multiple-gpu training script. We evaluate TGL using their single-gpu training script which does not have the issue and results in the worse prediction for MAG as shown in Table 4.

### v2

@@ Appendix D Dataset Description
  • GDELT [^4] https://github.com/tedzhouhk/TGL [25] is a Temporal Knowledge Graph (TKG) originated from the GDELT Event Database [52] and adapted by TGL for richer node and edge features. It is a large-scale dataset with more than 100M links. It records events from news and articles in over 100 languages every 15 minutes. The nodes represents actors and the edges represent events that happen between a pair of actors. The node features represent the CAMEO codes of the actors and the link features represent the CAMEO codes of the events. The labels are the countries where the actors were located when the events happen. According to the Terms of Use in the website (https://www.gdeltproject.org/), it is an open platform for research and analysis of global society and thus all datasets released by the GDELT Project are available for unlimited and unrestricted use for any academic, commercial, or governmental use of any kind without fee. The derived dataset by Zhou et al. [25] is open-sourced here.
@@
    MAG [^5] https://github.com/tedzhouhk/TGL Zhou et al. [25] is a homogeneous sub-graph of the heterogeneous MAG240M graph in OGB-LSC [53] extracted by TGL. It is another large-scale dataset with more than 100 million nodes and 1.3 billion links. It is a paper-paper citation network where each node in MAG represents one academic paper. The temporal edges represent citation of one paper to another with timestamp representing the year when the paper is published. The node features are the embeddings of the abstract of the paper generated with RoBERTa [54]. The node labels are the arXiv subject areas. The original dataset is licensed ODC-BY License. The derived dataset by Zhou et al. [25] is open-sourced here.
@@
  • Ubuntu [^6] https://snap.stanford.edu/data/sx-askubuntu.html [55, 56] is a dataset that records the events on a stack exchange web site called Ask Ubuntu. [^7] http://askubuntu.com/ The nodes represent users and there are three different types of edges, (1) user $u$ answering user $v$’s question, (2) user $u$ commenting on user $v$’s question, and (3) user $w$ commenting on user $u$’s answer. The original data dump is cc-by-sa 4.0 licensed. The derived dataset by Paranjape et al. [56] is open-sourced here.
@@
  • Wiki-talk [^8] https://snap.stanford.edu/data/wiki-talk-temporal.html [57, 56] is a dataset that represents the edit events on Wikipedia user talk pages. The dataset spans approximately 5 years so it accumulates a large number of nodes and edges. This is a large-scale dataset with more than 1M nodes. The original data dump is under the Creative Commons Attribution-Share-Alike 3.0 License. The derived dataset by Paranjape et al. [56] is open-sourced here.
@@
Table 9: Number of sampled temporal neighbors for each dataset for all methods. GDELT and MAG use smaller $s$ because of CPU and GPU size limits.
Params Wikipedia Reddit GDELT MAG Ubuntu Wiki-talk
@@
Table 10: The batch sizes used for scalability evaluation for each dataset for all methods.
@@ Appendix E Baselines, hyperparameters and the experiment setup
TGAT [17] adapts GAT [58] for static graphs to dynamic graphs. It aggregates messages from dynamic neighbors via an attention mechanism. While it proposed recent sampling, its implementation is significantly inefficient because it has to calculate the recency weights online before sampling. Thus, we only consider uniform sampling and truncation. We use 2 attention heads and 100 hidden dimensions. We only consider aggregation from the first-hop neighbors because aggregating second-hop neighbors introduces even more severe latency for sampling. We also do not observe significant improvement in prediction performance using second-hop neighbors. We do not find an official licensed original implementation for this work and we adopt the implementation by TGL [25].
@@
TGN [15] is a recently proposed SOTA temporal graph representation learning approach. It keeps track of a memory state for each node and update with new interactions. We train TGN with 100 dimensions for each of memory module, time feature and node embedding. We also only consider sampling the first-hop neighbors because of computational efficiency. Their source code [^9] https://github.com/twitter-research/tgn is licensed under the Apache-2.0 License but we adopt the implementation by TGL.
@@
APAN [21] is a recent approach designed for low inference latency. It prepares the neighborhood information forward with graph propagation instead of aggregation. While it inspires our approach, it does not support recent sampling and has shown to consume significant CPU memory while performing subpar to SOTA methods. The reason is it stores the messages for each node in CPU which consumes massive CPU memory. We train APAN with 100 dimensions for the node embeddings and 10 dimensions for the mailbox. Their original implementation [^10] https://github.com/WangXuhongCN/APAN is licensed under the MIT License but we adopt the implementation by TGL.
@@
TGL [25] is a very recent framework that support efficient and scalable temporal graph representation learning. It proposes temporal neighbor sampling with multi-core CPUs in parallel. Overall, it is shown to achieve on average 13$\times$ speedup for training and 173$\times$ speedup on neighbor sampling. For all of the baselines above, we evaluate them using the implementation by TGL. For the GDELT and the MAG datasets, since the node and edge features do not fit in our GPU, we set the all_on_gpu flag to false. For all other datasets, we set it to true. This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. The source code licensed under the Apache-2.0 License is provided here. [^11] https://github.com/tedzhouhk/TGL
@@
NAT [18] is also a very recent work that achieves SOTA link prediction performance for dynamic graphs. It proposed dictionary-type node representations to replace the traditional single-vector node representations which keep track of historical neighborhood information and avoid the expensive backward neighbor sampling. However, it relies on the construction of joint neighborhood features which only work with node pairs and cannot be easily generalized to node classification. Therefore, we evaluate it on link prediction only. We adapted NAT for node classification (called NAT-node) by removing the joint structural features and only aggregating their neighborhood representations to generate node representations. We also only use the first-hop neighborhood. Other than that, we follow the hyperparameter settings documented in their paper. Their implementation [^12] https://github.com/Graph-COM/Neighborhood-Aware-Temporal-Network is licensed under the MIT License.
@@
NLB (our work). For both NLB-node and NLB-edge, we tune the node representation and the node status dimensions between 50 to 100. We use two attention heads and we tune $\alpha$ and dropout between 0 to 1 with grid search.
Lastly, we note that the Adam optimizer [59] is used for all baselines.
@@ E.1 Inductive evaluation of NLB
Our evaluation pipeline for inductive learning is similar to NAT [18] and different from other baselines. For backward sampling methods such as TGN [15] and TGAT [17], when they do inductive evaluations, they can access the entire training and evaluation data for aggregation, including events that are masked for inductive test. However, NLB’s forward sampling prepares the down-sampled neighbors forward for future use and ignores the masked events during the training phase. Thus, by the end of the training, even all historical events become accessible, NLB cannot leverage them. Therefore, to ensure a fair comparison, after training, NLB processes the full train and validation data with all nodes unmasked to gather the down-sampled temporal neighbors from the complete set, and then processes the test data. Note that in this last pass over the full train and validation data, we only perform the forward sampling and do not perform training anymore.
@@ E.2 Node classification preprocessing
Appendix F Evaluation on TGB
TGB [44] is adopted by a lot of TGRL methods for standardizing the evaluation. Our main evaluation does not use TGB because it is insufficient for computational evaluation which involves billion-scale graphs. We provide some additional experiments for NLB using TGB here in Table 11. It verifies that NLB can match or outperform TGN in link prediction.
Method tgbl-review tgbl-coin tgbl-comment
TGN 34.90 $\pm$ 2.00 58.60 $\pm$ 3.70 37.90 $\pm$ 2.10
DyRep 22.00 $\pm$ 3.00 45.20 $\pm$ 4.60 28.90 $\pm$ 3.30
EdgeBank(tw) 2.50 58.00 14.90
NLB-edge 35.08 $\pm$ 2.20 61.92 $\pm$ 1.43 38.84 $\pm$ 2.17
NLB-node 37.13 $\pm$ 0.40 58.66 $\pm$ 2.14 36.91 $\pm$ 0.71
Table 11: Link prediction performance in Mean Reciprocal Rank (MRR) evaluated on TGB.
Appendix G TGL Node Classification for MAG
We fail to reproduce the node classification scores for MAG over all of the baselines implemented in TGL [25]. Our hypothesis is as follows. TGL has reported the scores for MAG based on an implementation with an issue. For the backtracking and sampling of temporal neighbors to learn representation at time $t$, TGL considers temporal neighbors of the same timestamp $t$ which may cause information leakage. We noticed this issue in their multiple-gpu training script. We evaluate TGL using their single-gpu training script which does not have the issue and results in the worse prediction for MAG as shown in Table 3.
@@ Appendix G TGL Node Classification for MAG
Appendix H The large difference between CPU and GPU energy consumption
While we directly obtained the numbers from the PyJoules [^13] https://pyjoules.readthedocs.io/en/latest/ output, we observe a large difference between CPU and GPU energy consumption. We find that it can be attributed to the following factors:
(1) Both the CPU and GPU have default energy costs even when no applications are running, as there may be background applications that constantly require energy. Within 1 minute, the default energy cost for the CPU is roughly 9,000 MJ, while the GPU cost is  0.8MJ.
(2) The training process is notably more energy-intensive on the CPU compared to the GPU, as there can be a considerable amount of CPU-GPU communication that is also counted on the CPU cost side as CPU controls the communication. This difference can be even more significant for applications that sample on the CPU side. For the NLB, we observed an additional  2,500MJ CPU energy and  12MJ GPU energy for training on the Wikitalk dataset. For TGN-uniform, it requires  10, [^0] MJ CPU energy and  10MJ GPU energy in extra.
Appendix I The efficiency evaluation of different sampling methods
In this section, we give additional details to show that forward sampling is indeed more efficient than samplings used by the conventional TGRL methods in practice. During the inference phase, both the proposed forward sampling and truncation methods can retrieve the down-sampled neighbors of a node in $O(1)$ time, while uniform sampling takes $O(N_{u}^{t})$ time, where $N_{u}^{t}$ is the number of neighbors of node u at time t. Since we typically need to infer a batch of nodes at a time, we prefer the sampling for the entire batch to be conducted in parallel. Both truncation and uniform sampling require storing and sampling from historical interactions. Due to the uneven distribution of degrees among different nodes, these operations can only be carried out within the CPUs, even when performed in parallel. On the contrary, our method can directly retrieve the already sampled neighbors within the GPU in constant time.
We have conducted an experiment that measures only the accumulated sampling time during inference for one epoch in seconds for three datasets (Table 12). The results show that our method can retrieve the samples instantaneously, while both truncation and uniform sampling are significantly slower in comparison. Uniform sampling is even slower than truncation, and multi-thread (32 threads) sampling is faster than single-thread sampling, especially for large datasets, as expected.
dataset proposed sampling trunc. 1-thread trunc. 32-thread unif. 1-thread unif. 32-thread
Wiki-talk 0.19 6.06 5.47 8.90 6.53
Reddit 0.10 0.96 0.68 1.02 0.82
GDELT 0.55 119.19 93.46 390.24 181.47
Table 12: The accumulated sampling time for inference during the training process over one epoch in seconds.
task dataset 0.2 0.4 0.6 0.8 0.9 1.0
NLB-edge trans. Reddit $72.39\pm 2.15$ $72.86\pm 2.06$ $73.13\pm 3.44$ $73.47\pm 2.48$ $75.40\pm 1.85$ $72.77\pm 2.69$
NLB-node $66.21\pm 2.95$ $64.85\pm 3.83$ $67.27\pm 2.00$ $68.15\pm 1.45$ $67.65\pm 5.31$ $67.85\pm 2.72$
NLB-edge induc. Ubuntu $21.59\pm 4.83$ $21.74\pm 2.84$ $22.72\pm 3.35$ $23.84\pm 2.27$ $21.13\pm 2.44$ $21.49\pm 4.49$
NLB-node $22.73\pm 3.07$ $21.08\pm 3.01$ $20.60\pm 4.22$ $20.54\pm 3.48$ $21.46\pm 2.97$ $21.30\pm 4.92$
Table 13: The performance measured in MRR with 500 negative samples of NLB-edge and NLB-node given different $\alpha$’s. $\alpha=0.9$ is the setting used for comparison with baselines.
task 0 5 10 15 20 25 30
NLB-edge trans. $62.35\pm 1.30$ $70.62\pm 3.16$ $69.93\pm 4.06$ $73.13\pm 1.98$ $75.40\pm 1.85$ $73.52\pm 1.34$ $72.30\pm 3.38$
NLB-node $62.35\pm 1.30$ $66.14\pm 4.74$ $68.96\pm 1.67$ $69.17\pm 1.82$ $67.65\pm 5.31$ $67.10\pm 4.42$ $66.78\pm 1.94$
NLB-edge induc. $36.59\pm 3.49$ $54.65\pm 4.36$ $59.62\pm 1.03$ $56.09\pm 5.45$ $61.74\pm 1.85$ $61.87\pm 2.03$ $61.33\pm 2.65$
NLB-node $36.59\pm 3.49$ $47.69\pm 6.44$ $51.15\pm 5.36$ $54.50\pm 4.76$ $53.43\pm 8.03$ $57.06\pm 7.02$ $58.09\pm 0.93$
Table 14: The performance measured in MRR with 500 negative samples of NLB-edge and NLB-node on the Reddit dataset given different down-sampled neighbor hash table sizes $s$. $s=20$ is the setting used for comparison with baselines.